OpenStudy (anonymous):

Simplify 3 square root of 125

OpenStudy (anonymous):

\[3\sqrt{125}\] does it look like this?

OpenStudy (anonymous):

Yes it does

OpenStudy (anonymous):

Can you show me how to do it please @jonnyrules96

OpenStudy (anonymous):

ok so first you want to take 125 and see what it is divisible by

OpenStudy (anonymous):

and how does one do that

OpenStudy (anonymous):

i found that it is divisible by 25 so since you know that you can make the equation look a little bit like this \[3\sqrt{25*5}\] now it just so happens that 5 is the square root of 25 so we can then make the equation look like this \[3*5\sqrt{5}\] and so completely simplified it should look like \[15\sqrt{5}\]

OpenStudy (anonymous):

Thank You Thank You Thank You Thank You

OpenStudy (anonymous):

no problemo =D

OpenStudy (anonymous):

wait how does one do a question like this where theres an exponent \[3\sqrt[4]{112}\]

OpenStudy (anonymous):

the exponent in this case just means that the 112 is the 4th power root rather then the square root (not sure if thats the terminology for it) so lets say you want to bring the 3 back into the original 4th power number then it would be \[3^{4}\] so just follow the same process as before except instead of square root its the root of the 4th power

OpenStudy (anonymous):

i hope that makes sense :P

OpenStudy (anonymous):

yes it does thank you

OpenStudy (anonymous):

=D g'luck!